OpenStudy (anonymous):

Solve: 4y2 + 12y + 9 = 0

OpenStudy (anonymous):

hint... \[\LARGE x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] substitute.. \[\LARGE a=4 \quad \quad b=12 \quad \quad c=9\]